OpenStudy (loser66):

Solve near point 0 x^2y" + y =0 I got characteristic equation r^2-2+1 =0 and 2 complex roots are r = \(\frac{1}{2}\pm \frac{\sqrt{3}}{2}i\) but I don't know how to construct the series. Please, help

OpenStudy (loser66):

Here what I went so far check regular singular point: \[lim_{x\rightarrow 0}x^2\frac{1}{x^2}=1 \] so, x =0 is regular singular point. and the stuff above.

OpenStudy (ybarrap):

This is an Euler's differential equation. You can solve using http://en.wikipedia.org/wiki/Euler%27s_differential_equation . Solution will involve x^r

OpenStudy (ybarrap):

You can probably use the limit idea or linearize around zero using taylor expansion

OpenStudy (loser66):

thanks a lot. I will.

OpenStudy (anonymous):

You shouldn't be getting a characteristic equation here...

OpenStudy (loser66):

why? that 's what I read from the book

OpenStudy (loser66):

is it Euler's form?

OpenStudy (anonymous):

No, the characteristic equation (only) arises from an equation that is linear, homogeneous, and has constant coefficients.

OpenStudy (loser66):

if you said so, what should I do? construct as y = sum a_n x^n?

OpenStudy (anonymous):

That's what I would do, given that the instruction say "Solve near point 0." Otherwise, @ybarrap is right about the \(y=x^r\).

OpenStudy (loser66):

Cannot be \[y =\sum_{n=o}^{infty} a_nx^n\]because after putting x^2y" +y =0 I get nothing.

OpenStudy (loser66):

Let sumation aside, y = a_n x^n y'= n a_n x ^(n-1) y"= n (n-1) a_n x ^(n-2) and then x^2y"= n(n-1) a_n x^n what can I get when putting them into the original equation?

OpenStudy (anonymous):

Hmm, in that case stick with ybarrap's suggestion.

OpenStudy (loser66):

so? don't have series solution, right?

OpenStudy (anonymous):

Yes, it would seem so...

OpenStudy (loser66):

Ok, thanks

OpenStudy (loser66):

@Microrobot

OpenStudy (anonymous):

Hi Loser, how can I help?

OpenStudy (loser66):

what do you mean by "how can"? you don't know this stuff?

OpenStudy (loser66):

I have a question: 1.2 E -7 means?

OpenStudy (ybarrap):

You still working this problem?

OpenStudy (ybarrap):

.00000012

OpenStudy (ybarrap):

Take the decimal point 7 units to the left, when minus, to the right if plus

OpenStudy (loser66):

yes, many other thing. and another question. why those are the same? please help me figure out

OpenStudy (ybarrap):

you mean, scientific notation vs. decimal format?

OpenStudy (loser66):

1) \[38*10^6 e^{-4*10^7 t}-38*10^6e^{-20*10^5t}+12*10^{-8}\] t = 0.001 and t = 0.01 I got the same answer , ha!! I know it's wrong but why and how to get right?

OpenStudy (ybarrap):

So with something like this it's very easy to make a calculator mistake. You know how to compute it, just need to go slow sometimes. Let's take this apart.

OpenStudy (ybarrap):

First term: $$\Large \bf 38*10^6e^{-4*10^7t}$$ Plug it t=.001 and we'll compare. Let me know when you are done.

OpenStudy (loser66):

0

OpenStudy (loser66):

how can it be, right? but I did it twice, still get 0

OpenStudy (ybarrap):

I get for t=.001: $$\bf \Large 6.31695... × 10^{-17365}$$ What do you get for \( \bf 38*10^6\)?

OpenStudy (loser66):

38 and 6 zero after it.

OpenStudy (ybarrap):

I think I see the problem. Your calculator can not handle large numbers. Try google: https://www.google.com/#bav=on.2,or.r_cp.r_qf.&fp=584e6f03ba966077&q=38*10%5E6

OpenStudy (ybarrap):

Google's not any better. Try wolfram: http://www.wolframalpha.com/input/?i=38*10%5E6e%5E%28-4*10%5E7%28.001%29%29

OpenStudy (loser66):

yes, the same me 1.2 *10^-7

OpenStudy (ybarrap):

What is this for? The total or just for the 1st term?

OpenStudy (loser66):

I don't know what to do next, can you help me check my work? especially at \(\lambda_ 2\) from my calculator, it showed me -20. 00008. but I really don't know how much it is, I assume that it is -20*10^5 , or it's just -20 ?? shame on me, differential equation student doesn't know how to read a number!!!

OpenStudy (ybarrap):

For \(\lambda_2\) I get -20.

OpenStudy (loser66):

Ok, thanks, I will redo my work. Ha, it leads to the wrong answer for the rest part, right??

OpenStudy (loser66):

so, lambda 1 is correct, right?

OpenStudy (ybarrap):

y(t) sort of cut off, but this is the final solution along with steady state: $$ \small y(t)=\frac{3 \left(\frac{250}{\sqrt{62499}}-1\right) \left(\left(124999+500 \sqrt{62499}\right) \left(-e^{-\frac{10000 t}{250+\sqrt{62499}}}\right)+e^{-10000 \left(250+\sqrt{62499}\right) t}+500 \sqrt{62499}+124998\right)}{50000000} \\\\ $$ Steady-state: $$ \bf \lim_{t\to\infty}y(t)=\frac{3 \left(\frac{250}{\sqrt{62499}}-1\right)(500 \sqrt{62499}+124998)}{50000000} $$ Details: http://www.wolframalpha.com/input/?i=%28-5*10%5E6-sqrt%285*10%5E12-4*10%5E8%29%29%2F2

OpenStudy (anonymous):

Im so sorry Loser. I was afk for some time.

OpenStudy (anonymous):

I meant by how can I help you is what you need.

OpenStudy (anonymous):

Sorry about that :(

OpenStudy (anonymous):

@ybarrap Thank you so much for helping him while I was gone!

OpenStudy (ybarrap):

ur welcome